Find the number of positive integers less than 601 that are not divisible by 4 or 5 or 6


1
Expert's answer
2021-11-08T08:09:16-0500

The number of positive integers less than 601 that are divisible by 4 


"\lbrack\dfrac{600}{4}\rbrack=150"

The number of positive integers less than 601 that are divisible by 6 


"\lbrack\dfrac{600}{6}\rbrack=100"

The number of positive integers less than 601 that are divisible by 5 


"\lbrack\dfrac{600}{5}\rbrack=120"

The number of positive integers less than 601 that are divisible by 4 and 6


"\lbrack\dfrac{600}{12}\rbrack=50"

The number of positive integers less than 601 that are divisible by 4 and 5 


"\lbrack\dfrac{600}{20}\rbrack=30"

The number of positive integers less than 601 that are divisible by 5 and 6 


"\lbrack\dfrac{600}{30}\rbrack=20"

The number of positive integers less than 601 that are divisible by 4, and 5, and 6 


"\lbrack\dfrac{600}{60}\rbrack=10"

The number of positive integers less than 601 that are divisible by 4 or 5 or 6


"150+100+120-50-30-20+10=280"

The number of positive integers less than 601 that are not divisible by 4 or 5 or 6


"600-280=320"